OpenStudy (anonymous):

please help me with this question! Using the completing-the-square method, rewrite f(x) = x2 − 8x + 3 in vertex form. f(x) = (x − 8)2 f(x) = (x − 4)2 − 13 f(x) = (x − 4)2 + 3 f(x) = (x − 4)2 + 16

OpenStudy (anonymous):

The question is Using the completing-the-square method, rewrite f(x) = x2 − 8x + 3 in vertex form.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi!

OpenStudy (misty1212):

do you have to show your work or can you just do it?

OpenStudy (anonymous):

you don't have to show your work

OpenStudy (misty1212):

ok then lets to it quick you have \[f(x)=x^2-8x+3\] right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

what is half of 8?

OpenStudy (anonymous):

4

OpenStudy (misty1212):

so go right to \[f(x)=(x-4)^2+k\]

OpenStudy (misty1212):

to find \(k\) plug in 4 in to \[f(x)=x^2-8x+3\]

OpenStudy (anonymous):

so I solve f(x)= (x-4)^2 +k and for k, I plug in 4?

OpenStudy (misty1212):

yes, let me know what you get and i will check it

OpenStudy (anonymous):

okay I'll solve it real quick :)

OpenStudy (anonymous):

I squared the -4 in the parenthesis and got f(x)= (x-16)+4

OpenStudy (anonymous):

is that all that I have to do?

OpenStudy (misty1212):

whoa hold on

OpenStudy (misty1212):

lets back up a second

OpenStudy (misty1212):

we got half of 8 is 4, so it s going to look like \[f(x)=(x-4)^2+k\]

OpenStudy (misty1212):

now we need the number \(k\) right?

OpenStudy (anonymous):

yes and you said to plug 4 in for k right?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

to find \(k\) start with \[f(x)=x^2-8x+3\] and find \(f(4)\) \[f(4)=4^2-8\times 4+3\]

OpenStudy (misty1212):

let me know what you get now

OpenStudy (anonymous):

oh! okay. I totally switched up what you said :o I'll re-solve it

OpenStudy (anonymous):

I got -13 on my calculator, but when I solved it on paper I got 35. Sorry if I'm annoying you, I just don't understand algebra 2 at all :/

OpenStudy (misty1212):

it is ok dear, not problem the correct one is \(-13\) lets do it with pencil and paper

OpenStudy (misty1212):

\[f(4)=4^2-8\times 4+3\] first square \[16-8\times 4+3\] then multiply \[16-32+3\] then subtract \[-16+3\] then add \[-13\]

OpenStudy (misty1212):

final answer to your question "vertex form" is \[f(x)=(x-4)^2-13\]

OpenStudy (anonymous):

Oh! I solved the exponent first and then subtracted. But thank you so much for helping me! :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Hey @misty1212 . Can you help me with one more question? It also involves the vertex form and standard form

OpenStudy (misty1212):

sure happy to!!

OpenStudy (anonymous):

It's this - Rewrite f(x) = –2(x − 3)2 + 2 from vertex form to standard form.

OpenStudy (misty1212):

ok this one is going the other way around from vertex to standard how is your algebra?

OpenStudy (misty1212):

because that is what you need \[-2(x-3)^2+2=-2(x-3)(x-3)+2\] you have to multiply out, distribute, combine like terms

OpenStudy (anonymous):

Should I solve that and then tell you what I got?

OpenStudy (misty1212):

sure don't forgot to multiply out first \[(x-3)(x-3)\] requires 4 multiplications then multiply each term by \(-2\)

OpenStudy (anonymous):

So for the first part of the equation, I got -2x+6^2 +2. Is that right so far?

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

\(-3\times (-3)=+9\) lets go slow

OpenStudy (misty1212):

\[(x-3)(x-3)\]first we get \[x^2-3x-3x+9\] let me know when that is clear, or if not, ask

OpenStudy (anonymous):

Yes that makes sense

OpenStudy (misty1212):

you sure?

OpenStudy (misty1212):

you see where each term came from?

OpenStudy (misty1212):

some times math teachers call it "foil"

OpenStudy (anonymous):

Yeah. The x's get multiplied to each other and that makes x^2, and then you multiply the 3's to the x's and get -3x-3x. Last, you multiply the -3's to each other and you get 9. But how would we put that into standard form?

OpenStudy (anonymous):

My answers are f(x) = –2x2 − 18 f(x) = –2x2 + 12x + 20 f(x) = –2x2 + 12x − 16 f(x) = 4x2 − 24x + 38

OpenStudy (misty1212):

ok so we are at \[x^2-3x-3x+9\] which is the same as \[x^2-6x+9\]

OpenStudy (misty1212):

then \[-2(x^2-6x+9)=-2x^2+12x-18\]

OpenStudy (misty1212):

then add the \(2\) at the end, get \[x^2+12x-16\]

OpenStudy (anonymous):

So the answer would be f(x)= -2x^2+12x-16?

OpenStudy (misty1212):

yes FLVS?

OpenStudy (anonymous):

yes. Do you do FLVS?

OpenStudy (misty1212):

lol no dear, i don't

OpenStudy (anonymous):

oh haha

OpenStudy (misty1212):

good luck with the algebra \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Thank you so much! And thanks for the help again :)